OpenStudy (anonymous):

Verify that the Intermediate Value Theorem applies to the indicated interval and find the value of "c" guaranteed by the theorem. f(x)=(x^2) + x - 1, [0,5] f(c) =11

OpenStudy (anonymous):

do you just plug in 11?

OpenStudy (anonymous):

well you first show that 11 is between f(0) and f(5) then in order to find c you solve: 11 = c^2 +c - 1

OpenStudy (anonymous):

how do you show that 11 is between f(0) and f(5)?

OpenStudy (anonymous):

what is f(0) ?

OpenStudy (anonymous):

do you just plug in 0 to the equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

f(0) = -1

OpenStudy (anonymous):

good and f(5) ?

OpenStudy (anonymous):

29

OpenStudy (anonymous):

so 11 is really between them and according to the theorem there must be a value for c such that f(c) = 11

OpenStudy (anonymous):

now find c 11=c^2 + c - 1

OpenStudy (anonymous):

do I add over the 1 or subtract the 11?

OpenStudy (anonymous):

@Coolsector

OpenStudy (anonymous):

you better subtract the 11 so you will get quadratic equation

OpenStudy (anonymous):

(x+4)(x-3)=0

OpenStudy (anonymous):

but the answer in my book says f(3)=11, why not use f(-4) ?

OpenStudy (anonymous):

because the interval was [0,5]

OpenStudy (anonymous):

OH. wow okay LOL thank you

OpenStudy (anonymous):

yw:)